OpenStudy (anonymous):

Find the value of c that is guaranteed by the Mean Value Theorem for the function f(x)= x^3-12x on the interval [-3,0].

zepdrix (zepdrix):

So the `Mean Value Theorem` tells us: Given a `secant line`, (we're using the end points of our interval for our secant line), If the function is continuous, there MUST exist a `tangent line` somewhere in the interval that has the same slope as that secant line. So let's first find the value of the secant line. \[\large \frac{f(0)-f(-3)}{0-(-3)}\] So here is the slope of our secant line. Let's plug in some values and figure out what that slope is.

zepdrix (zepdrix):

f(0)=0. Plugging in the -3's for f(-3) gives us something like this, \[\large \frac{0-\left[(-3)^3-12(-3)\right]}{3}\]

zepdrix (zepdrix):

Understand that part? :O

OpenStudy (anonymous):

yup!

zepdrix (zepdrix):

So what do you get for the slope of the secant line?

zepdrix (zepdrix):

-3 I think? When you simplify it all down. I dunno, I was trying to do the math in my head lol.

OpenStudy (anonymous):

i got the same answer as yours, but the answers that the book is giving is different. a) -sqrt(3) b) sqrt(3) c) -2sqrt(3) d) -sqrt(5) e)sqrt(5)

zepdrix (zepdrix):

Yes we still have a bit of work to do :)

zepdrix (zepdrix):

So we've found the slope of the secant line. The Mean Value Theorem is telling us that there exists a tangent line with the same slope value. In other words,\[\large f'(c)=-3\] Where \(\large c\) is somewhere inside of our interval. That's the value we want to solve for. The derivative at a particular point represents the slope of a `tangent line`. So this will give us what we want. We'll first want to take the derivative of our function f.\[\large f(x)=x^3-12x\]\[\large f'(x)=?\]

OpenStudy (anonymous):

3x^2-12

zepdrix (zepdrix):

Ok good. From here, we'll plug in a c for our x. (This is necessary, it just follows the notation of the Theorem).\[\large f'(c)=3c^2-12\] Since we already determined that \(\large f'(c)=-3\) We'll set these equal to one another,\[\large 3c^2-12=-3\] And solve for c! If you get 2 values for c, think about which one falls inside of our interval.

zepdrix (zepdrix):

This is not necessary* Blah typo

OpenStudy (anonymous):

so the correct answer will be sqrt(3) :)

zepdrix (zepdrix):

Ummm let's see. Dividing everything by 3,\[\large c^2-4=-1\] Doing some math...\[\large c=\pm \sqrt3\] We have to remember a positive and negative solution come out taking the root of a square. Which one is the c value that we want? Which one falls in our interval, [-3,0]?

OpenStudy (anonymous):

sqrt(3)

OpenStudy (anonymous):

no, -sqrt(3)

zepdrix (zepdrix):

Yah the negative one falls in our interval right? Good job.